Give the equation of a line perpendicular to y=2x+4 that passes through the point (-8,3) and identify the point of intersection as an ordered pair.
I really need this for my final exam and i have to graph it so if i can get the explanation and answer it would be lovely thank you very much.

Answer by MsRedz123(24)

In y=2x + 4 the slope is 2. So, it's slope perpendicularly is (-1/2).
m=(-1/2) (-8,3)
Y-Y1=m(X-X1)
Y-3=-1/2(x+8)
Y-3=-x/2-4
2[(y-3)=-x/2-4]
2Y-6=-x-8
x+2y-6+8=0
x+2y+2=0
Then, y=2x+4
x+2y+2=0
Solve for x using y.
x+2(2x+4)+2=0
x+4x+8+2=0
5x+10=0
5x=-10
x=-2
Solve for y using the value of x.
y=2x+4
y=2(-2)+4
y=-4+4
y=0
The intersection therefore is (-2,0)